ULUS 8.53. but include in the s an isentropic efficiency 8.54 Reconsider the cycle o f lem 852 L analysis that each turbine e8.55
Answer #1

geansabay Steam (4) (8) tus Conolensoy OFWH. state - 1 Palesure (P.) = 16 MPa. Temperature (Ti) = 560°C State - Super headedOhMos pon ding to Preesule of 4MAX. Enthalpy cat lule 2 le Enthalpy and Enbopy = 8 = 6. 5132 x/leg the = 3050.86 KT/Reg stateNow, Enthalpy (hz) = het dalhg-he) hz = 561.4 +0.9112725. 3- 561.457) hz = 2530.48 ks) Reg. Sinuce (se LS4L Sg ) at 8 kPa so,Now at Eupa Enthalpy of całwiated viapouse (he) = 2549.0 Wilforen Enthalpy welcsaturated liquid the) = 173-88 KI/kg Enthalpystate- 8 Pressure (Pg) = 16 MPa Enthalpy (he) hot + V ( Prz - P6) he 561.477 + (1.0732x103) (16 - 0.3] x 103. hg = 561.47 + 1state-1 Since Espansion in tropis luo enthalpie in nature so, Enthalpy cat estate 41 is some cas og Enthalpy of whate 10 hu =ay Net Power developed - EWT - Ewp. Whet miskam [pl hp the) +(1+y) (h2-hg) + (1-41-42) Che-huly - {(1-4=Y2) ( ho-hy) + (he-hqTaking Specific heat is coding coater C = 4.187 kilka c. The Temperature difference of water (AT) = 18°C Let me be the minus